# SpikeNAS: A Fast Memory-Aware Neural Architecture Search Framework for Spiking Neural Network-based Autonomous Agents

### v3

@@ III-C A Fast Memory-aware Search Algorithm
  • Each cell has an individual search to determine its architecture, thereby exploring more candidates from different combinations of architectures across multiple cells (Alg. 1: line 2).
  • Investigation of redundant architectures for each cell is minimized to improve the effectiveness of our architecture search. To do this, we propose an in-cell architecture search (Alg. 1: line 3-14). Variables $con_{xy}$ with $xy$$\in\{01,02,03,12,13,23\}$ in Alg. 1: line 9-14 correspond to the type of operation that connect node-$X$ to node-$Y$; see Fig. 7. Then, the cell architecture is constructed once $con_{xy}$ conections are defined (Alg. 1: line 15).
  • All cells are initiated with the same architecture (Alg. 1: line 16-17). Then, the architecture candidates are explored for each cell, starting from the second cell to the last one, consecutively (Alg. 1: line 18-19). Hence, the search space is reduced while keeping good architectures from previous cells. Then, the network architecture is constructed once all cells are defined (Alg. 1: line 21).
  • Information of the memory constraint is included into the search process (Alg. 1: line 22). It is used to select which architectures that should be evaluated, hence filtering out the ones that are not suitable and optimizing the searching time even further (Alg. 1: line 23-27). To enable this, an analytical model of memory footprint is developed and employed into the search. Details of the proposed analytical model is discussed in the following.
Algorithm 1 Our Fast Memory-aware Search Algorithm
0:  (1) Pre-defined number of neural cells ($C$); (2) Neural cells: object names ($cell_{1}$, …, $cell_{C}$), neural cell architecture ($arch$) ;(3) Memory constraint or budget ($mem_{budget}$);(4) Pre-defined number of operation types ($T$); (5) Operation codes: SkipCon (0), 3x3Conv (1), and 3x3AvgPool (2);
0:  Generated SNN architecture ($net\_arch$); BEGIN Initialization:
1:  $score_{h}$=-1000; Process:
2:  for ($i$ = 1; $i$¡($C$+1); $i$++) do
3:     for ($c01$=0; $c01$¡$T$; $c01$++) do
4:        for ($c02$=0; $c02$¡$T$; $c02$++) do
5:           for ($c03$=0; $c03$¡$T$; $c03$++) do
6:              for ($c12$=0; $c12$¡$T$; $c12$++) do
7:                 for ($c13$=0; $c13$¡$T$; $c13$++) do
8:                    for ($c23$=0; $c23$¡$T$; $c23$++) do
9:                       $con_{01}=c01$;
10:                       $con_{02}=c02$;
11:                       $con_{03}=c03$;
12:                       $con_{12}=c12$;
13:                       $con_{13}=c13$;
14:                       $con_{23}=c23$;
[…]
16:                       if ($i$==1) then
17:                          $cell_{1},...,cell_{N}\leftarrow arch$;
18:                       else
19:                          $cell_{i}\leftarrow arch$;
20:                       end if
21:                       $net$ = create$(cell_{1},...,cell_{C})$;
22:                       $mem_{cost}$= memcalc($net$);
23:                       if ($mem_{cost}\leq mem_{budget}$then
24:                          $score=$ eval$(net)$;
25:                          if ($score>score_{h}$then
26:                             $score_{h}=score$;
27:                             $net\_arch=net$;
28:                          end if
29:                       end if
30:                    end for
31:                 end for
32:              end for
33:           end for
34:        end for
35:     end for
36:  end for
37:  return  $net\_arch$; END
@@
$\begin{split}M_{param}=N_{param}\cdot bit\end{split}$ (8)
IV Evaluation Methodology
[…]
For evaluation and validation of the proposed concepts, we develop a Python-based implementation using the SpikingJelly library [40] on a multi-GPU machine (i.e., an Nvidia RTX 6000 Ada GPU machine with 48GB GDDR6 memory), and consider CIFAR10 and CIFAR100 as the workloads. We perform up to 300 training epochs with surrogate gradient learning, while evaluating the test accuracy gradually to see the learning curve of the investigated network architecture. These experiments provide information of accuracy, searching time, and memory footprint.
@@ IV Evaluation Methodology
TABLE I: Different scenarios for SpikeNAS with 3 operation types: SkipCon, 3x3Conv, and 3x3AvgPool.
@@
TABLE II: Different scenarios for SpikeNAS with 2 operation types: SkipCon and 3x3Conv.
@@
V Results and Discussion
Figure 11: Results of the accuracy for (a) CIFAR10 and (b) CIFAR100.
Figure 12: Results of the memory footprint normalized to the state-of-the-art work SNASNet_2C5O for (a) CIFAR10, and (b) CIFAR100.
Figure 13: Results of the searching time normalized to the state-of-the-art work SNASNet_2C5O for (a) CIFAR10, and (b) CIFAR100.
V-A Maintaining the High Accuracy
Fig. 11(a) and Fig. 11(b) present the experimental results for the accuracy for CIFAR10 and CIFAR100, respectively.
@@ V-A Maintaining the High Accuracy
The 2-Cell Architectures: SNASNet achieves 92.58% of accuracy for CIFAR10 and 69.18% for CIFAR100. Our SpikeNAS_2C3O achieves 93.06% of accuracy for CIFAR10 and 71.62% for CIFAR100, while SpikeNAS_2C2O achieves 93.4% of accuracy for CIFAR10 and 71.26% for CIFAR100; see 1 and 2. These results indicate that, our SpikeNAS improves the accuracy as compared to the state-of-the-art, since SpikeNAS only considers operations with high significance to the accuracy. Moreover, all of these accuracy improvements can be achieved by networks from SpikeNAS with less than 10M parameters. When the memory constraints are given, our SpikeNAS_2C3O_M achieves 90.96% of accuracy for CIFAR10 and 70.52% for CIFAR100, while SpikeNAS_2C2O_M achieves 92.12% of accuracy for CIFAR10 and 68.02% for CIFAR100; see 1 and 2. These results show that, our SpikeNAS maintains (or improves in some cases) the accuracy as compared to the state-of-the-art, while ensuring that the generated SNN architecture meets the memory budget. The reason is that, SpikeNAS employs operations with high significance to the accuracy and incorporates the memory budget in its search process. Note, memory evaluation is discussed in Sec. V-B.
@@
The 1-Cell and 3-Cell Architectures (Variants): Our SpikeNAS_1C3O achieves 91.29% of accuracy for CIFAR10 and 64.79% for CIFAR100, while SpikeNAS_1C2O achieves 91.60% of accuracy for CIFAR10 and 65.25% for CIFAR100; see 3 and 4. Furthermore, our SpikeNAS_3C3O achieves 93.51% of accuracy for CIFAR10 and 73.07% for CIFAR100, while SpikeNAS_3C2O achieves 93.43% of accuracy for CIFAR10 and 73.25% for CIFAR100; see 3 and 4. These results indicate that, in general, the 1-cell architectures offer a lower accuracy as compared to the 2-cell architectures, as the 1-cell architecture have fewer operations to extract meaningful features from the inputs. Meanwhile, 3-cell architectures can offer a higher accuracy as compared to the 2-cell architectures, as the 3-cell architecture have more operations to extract meaningful features from the inputs. When the memory constraints are given, SpikeNAS_1C3O_M and SpikeNAS_1C2O_M have the similar results to their unconstrained counterparts (i.e., SpikeNAS_1C3O and SpikeNAS_1C2O) as the memory footprints of the unconstrained ones are already within the given memory budgets. Meanwhile, SpikeNAS_3C3O_M and SpikeNAS_3C2O_M have lower accuracy than their unconstrained counterparts (i.e., SpikeNAS_3C3O and SpikeNAS_3C2O), since the memory constraint adds a limitation for finding a combination of cell architectures that offers high accuracy.
@@
V-B Ensuring Memory-aware SNN Generation
@@
Fig. 12(a) and Fig. 12(b) present the experimental results for the memory footprint for CIFAR10 and CIFAR100, respectively. For CIFAR10, SNASNet incurs 6% memory overhead than the given budget, while SpikeNAS can save the memory footprint up to 43% from the state-of-the-art; see 5. If we consider high accuracy, then we can select SpikeNAS_2C2O_M with 92.12% accuracy and 30% memory saving from the state-of-the-art (24% memory saving from the given budget); see 6. Meanwhile, for CIFAR100, SNASNet incurs 23% memory overhead than the given budget, while SpikeNAS can save the memory footprint up to 72% from the state-of-the-art; see 7. If we consider high accuracy, then we can select SpikeNAS_2C3O_M with 70.52% accuracy and 38% memory saving from the state-of-the-art (15% memory saving from the given budget); see 8. These results highlight that, SpikeNAS can generate SNN architectures that meet the given memory constraints due to its memory-aware search algorithm, enabling efficient deployments of SNN models on the target hardware platforms.
V-C Reducing the Searching Time
@@
Fig. 13(a) and Fig. 13(b) present the experimental results for the searching time for CIFAR10 and CIFAR100, respectively. In both cases of workloads (CIFAR10 and CIFAR100), our SpikeNAS offers significant improvements for the searching time as compared to the state-of-the-art. The reason is that, SpikeNAS employs fewer operation types that have significant impact on the accuracy, performs a consecutive search across the existing cells, and incorporates a memory constraint into the search process, thereby leading to a smaller search space and a faster searching time than the state-of-the-art. If we consider high accuracy, acceptable memory cost, and fast searching time, then we can select SpikeNAS_2C2O_M for CIFAR10 with 92.12% accuracy, 23% memory saving from the given memory budget, and 347.7x faster search than the state-of-the-art; see 9. For CIFAR100, we can select SpikeNAS_2C3O_M with 70.52% accuracy, 14% memory saving from the given memory budget, and 4.4x faster search than the state-of-the-art; see 10. These results highlight that, our SpikeNAS can quickly generate SNN architectures that offer high accuracy and meet the memory constraints, hence enabling efficient design automation for developing SNN-based autonomous mobile agents.
VI Conclusion

### v4

@@ 3.3 A Fast Memory-aware Search Algorithm
  • We employ a dynamic programming approach to perform architecture search. Specifically, we search and determine the architecture for each cell one-by-one (see Alg. 1: line 2). Therefore, more network candidates from different combinations of architectures across multiple cells are explored, hence leading to a higher possibility to find an SNN architecture that achieves high accuracy.
  • Investigation of redundant architectures for each cell is minimized to improve the effectiveness of our architecture search. To enable this, we propose an in-cell architecture search (as shown in Alg. 1: line 3). Here, variables $con_{xy}$ with $xy$$\in\{01,02,03,12,13,23\}$ correspond to the type of operation that connect node-$X$ to node-$Y$, as shown in Fig. 7. Then, the cell architecture is constructed once $con_{xy}$ connections are defined (see Alg. 1: line 4).
  • All cells are initiated with the same architecture (see Alg. 1: line 5-6). Then, the architecture candidates are explored for each cell, starting from the second cell to the last one, consecutively (see Alg. 1: line 7-8). Hence, the search space is reduced while keeping good architectures from previous cells. Then, the network architecture is constructed once all cells are defined (Alg. 1: line 10), and its memory cost is also evaluated (Alg. 1: line 11).
  • Afterward, the memory-constraining process is performed by incorporating the maximum allowed number of parameters into the search (see Alg. 1: line 12). It is used to select which architectures should be evaluated, hence filtering out the ones that are not suitable and optimizing the searching time even further. If the number of parameters from the investigated network ($mem_{cost}$) is less than or equal to the memory constraint/budget ($mem_{budget}$), then the network evaluation and selection are performed (see Alg. 1: line 13-16). To do this, an analytical model of memory footprint is developed and employed into the search. Details of the proposed analytical model is discussed in the following.
@@
$\begin{split}M_{param}=N_{param}\cdot bit\end{split}\vspace{-1cm}$ (8)
Algorithm 1 Our Fast Memory-aware Search Algorithm
0:  (1) Pre-defined number of neural cells ($C$); (2) Neural cells: object names ($cell_{1}$, …, $cell_{C}$), cell architecture ($arch$); (3) Memory constraint or budget ($mem_{budget}$); (4) Pre-defined number of operation types ($T$); (5) Operation ($P$): SkipCon=0, 3x3Conv=1, and 3x3AvgPool=2; hence $P\in\{0,1,2\}$
0:  Generated SNN architecture ($net\_arch$); BEGIN Initialization:
1:  $score_{h}$=-1000; Process:
2:  for ($i$=1; $i$$<$($C$+1); $i$++) do
3:     for each combination from ($\forall$ $con_{01}$in $P$), ($\forall$ $con_{02}$in $P$), ($\forall$ $con_{03}$in $P$), ($\forall$ $con_{12}$in $P$), ($\forall$ $con_{13}$in $P$), and ($\forall$ $con_{23}$in $P$do
4:        $arch$ = Cell($con_{01},con_{02},con_{03},con_{12},con_{13},con_{23}$); // create the cell architecture
5:        if ($i$==1) then
6:           $cell_{1},...,cell_{N}\leftarrow arch$; // apply $arch$ to all cells
7:        else
8:           $cell_{i}\leftarrow arch$; // apply $arch$ to $cell_{i}$
9:        end if
10:        $net$ = Net$(cell_{1},...,cell_{C})$; // create the network
11:        $mem_{cost}$= calc_mem($net$); // calculate the memory cost
12:        if ($mem_{cost}\leq mem_{budget}$then
13:           $score=$ eval$(net)$; // evaluate the network score
14:           if ($score>score_{h}$then
15:              $score_{h}=score$;
16:              $net\_arch=net$;
17:           end if
18:        end if
19:     end for
20:  end for
21:  return  $net\_arch$; END
3.4 SNN Quantization
To further reduce the memory footprint of generated SNNs, we quantize the SNN weights through a post-training quantization (PTQ) approach. To realize the quantization, we can employ the desired number format and rounding scheme. For instance, we consider the fixed-point format with truncation scheme, as they offer high accuracy while enabling fast design space exploration for SNN processing [26]. The fixed-point format can be represented in $b$-bit precision and follows the 2’s complement method, while the truncation (TR) scheme for the given number $x$ under $b$-bit precision can be represented as TR($x$,$b$) = $\lfloor x\rfloor$ [20]. However, quantizing the weight value also implies a reduction of its representation capability, thereby potentially decreasing the accuracy [17]. To address this, we consider the trade-off between accuracy and memory footprint to determine the acceptable quantization levels (i.e., bit precision). Therefore, we can select the acceptable accuracy and memory footprint to meet with the design specifications.
The above discussion shows that, the SpikeNAS framework aims to enable a fast memory-aware NAS for generating a suitable SNN based on the memory constraint (i.e., maximum allowed number of parameters) and the NAS without training approach, which has not been comprehensively studied by state-of-the-art works. To achieve this, our SpikeNAS performs new studies and employs novel optimization techniques that are tailored specifically for SNNs, including analysis of the impact of SNN operations, network architecture enhancements, and a fast memory-aware search algorithm. Moreover, we also enhanced the SpikeNAS framework with post-training quantization (PTQ) to further reduce the memory footprint of the generated SNN. In this manner, our SpikeNAS framework integrates new studies, novel optimization techniques, and a widely-used PTQ technique to achieve high accuracy and high energy efficiency SNN processing.
4 Evaluation Methodology
To evaluate our SpikeNAS framework, we deploy the experimental setup as presented in Fig. 10, while considering the same evaluation conditions as widely used in the SNN community (e.g., focusing on the feedforward topology) [49]. We evaluate different scenarios for SpikeNAS as presented in Table 1 and 2 to provide a comprehensive study. For brevity, we use an encoding for these scenarios: “SpikeNAS_pCqO” refers to an architecture with p cell(s) and q operation types without memory constraint, while “SpikeNAS_pCqO_M” refers to the one with memory constraint. We consider 1.2x106 (1.2M), 2x106 (2M), and 4x106 (4M) parameters as the memory constraints for CIFAR10, CIFAR100, and TinyImageNet200, respectively. For the comparison partner, we consider SNASNet [28] and its default settings for the search space: 2 cells, 5 pre-defined operation types, and 5000x search iterations using a random search for experiments (i.e., SNASNet_2C5O).
@@ 4 Evaluation Methodology
Table 1: Different scenarios for SpikeNAS with 3 operation types: SkipCon, 3x3Conv, and 3x3AvgPool.
@@
Table 2: Different scenarios for SpikeNAS with 2 operation types: SkipCon and 3x3Conv.
@@
Figure 11: Results of the accuracy for SNNs with (a) 2 cells, (b) variants of 1 & 3 cells, and (c) quantized weights across different NAS techniques, settings, and datasets: CIFAR10, CIFAR100, and TinyImageNet200.
For evaluation and validation of the proposed concepts, we develop a Python-based implementation using the SpikingJelly library [50] on the Nvidia RTX A6000 GPU machine with 48GB GDDR6 memory. We consider the CIFAR10, CIFAR100, and TinyImageNet200 datasets as workloads which represent simple, medium, and complex tasks, respectively. We perform up to 300 training epochs with the surrogate gradient learning [51], while evaluating the test accuracy gradually to see the learning curve of the investigated network architecture. These experiments provide information of accuracy, searching time, memory footprint, and power consumption. These searching time and power are then leveraged to estimate the energy consumption. Meanwhile, for evaluating the proposed concepts on embedded platforms, we employ the Python-based implementation on the Nvidia Jetson AGX Orin.
5 Results and Discussion
5.1 Maintaining High Accuracy
Fig. 11(a)-(b) present the accuracy for SNNs with 2 cells and variants (i.e., 1 and 3 cells), respectively. Meanwhile, Fig. 11(c) presents the accuracy for quantized SNNs.
@@ 5.1 Maintaining High Accuracy
The 2-Cell Architectures: SNASNet achieves 92.58% of accuracy for CIFAR10, 69.18% for CIFAR100, and 53% for TinyImageNet200. Our SpikeNAS_2C3O achieves 93.10% of accuracy for CIFAR10, 71.81% for CIFAR100, and 52% for TinyImageNet200; while SpikeNAS_2C2O achieves 93.05% of accuracy for CIFAR10, 71.34% for CIFAR100, and 63% for TinyImageNet200; see 1, 2, and 3. These results indicate that, our SpikeNAS maintains (or improves in most cases) the accuracy as compared to the state-of-the-art, since SpikeNAS only considers operations with high significance to the accuracy. Moreover, these improvements can be achieved by networks from SpikeNAS with less than 10M parameters. When the memory constraints are given, our SpikeNAS_2C3O_M achieves 91.13% of accuracy for CIFAR10, 67.06% for CIFAR100, and 72% for TinyImageNet200; while SpikeNAS_2C2O_M achieves 92.02% of accuracy for CIFAR10, 68.41% for CIFAR100, and 61% for TinyImageNet200; see 1, 2, and 3. These results show that, our SpikeNAS maintains (or improves in some cases) the accuracy as compared to the state-of-the-art, while ensuring that the generated SNN architecture meets the memory budget. The reason is that, SpikeNAS employs operations with high significance to the accuracy and incorporates the memory budget in its search process. Note, memory evaluation is discussed in Section 5.2.
Figure 12: (a) Results of the number of parameters normalized to the state-of-the-art work SNASNet_2C5O across different NAS techniques, settings, and datasets. (b) Accuracy-memory trade-offs between state-of-the-art work SNASNet_2C5O with our SpikeNAS_2C2O_M under different precision levels and across different datasets: CIFAR10, CIFAR100, and TinyImageNet200.
@@
The 1-Cell and 3-Cell Architectures (Variants): Our SpikeNAS_1C3O achieves 91.65% of accuracy for CIFAR10, 65.10% for CIFAR100, and 54% for TinyImageNet200; while SpikeNAS_1C2O achieves 91.41% of accuracy for CIFAR10, 64.92% for CIFAR100, and 51% for TinyImageNet200; see 4, 5, and 6. Furthermore, our SpikeNAS_3C3O achieves 93.16% of accuracy for CIFAR10, 73.74% for CIFAR100, and 62% for TinyImageNet200; while SpikeNAS_3C2O achieves 93.63% of accuracy for CIFAR10, 72.65% for CIFAR100, and 66% for TinyImageNet200; see 4, 5, and 6. These results indicate that, in general, the 1-cell architectures offer a slightly lower accuracy as compared to the 2-cell architectures, as the 1-cell architecture have fewer operations to extract meaningful features from the inputs. Meanwhile, 3-cell architectures can offer a higher accuracy as compared to the 2-cell architectures, as the 3-cell architecture have more operations to extract meaningful features from the inputs. When the memory constraints are given, SpikeNAS_1C3O_M and SpikeNAS_1C2O_M have similar results to their unconstrained counterparts (i.e., SpikeNAS_1C3O and SpikeNAS_1C2O) for both CIFAR10 and CIFAR100, since the memory footprints of the unconstrained ones are already within the given memory budgets, and hence leading to similar network architectures and accuracy. Meanwhile, SpikeNAS_3C3O_M and SpikeNAS_3C2O_M have lower accuracy than their unconstrained counterparts (i.e., SpikeNAS_3C3O and SpikeNAS_3C2O), since the memory constraint adds a limitation for finding a combination of cell architectures that offers high accuracy.
Accuracy under Quantization: In general, applying lower bit precision levels leads to accuracy degradation due to the reduced capabilities of representing more learned knowledge (information) in the weights. Here, we observe that employing at least 10-bit precision for CIFAR10 7 and CIFAR100 8 as well as at least 8-bit precision for TinyImageNet200 9 maintains the accuracy comparable to that of the state-of-the-art (and even improves the accuracy in some cases), indicating that these precision levels are sufficient for preserving the learned knowledge (i.e., input features).
@@
5.2 Ensuring Memory-aware SNN Generation
@@
Fig. 12(a) presents the experimental results for the number of parameters across different NAS techniques, settings, and datasets. For CIFAR10, SNASNet incurs 6% memory overhead than the given budget, while SpikeNAS can save the memory footprint up to 42% from the state-of-the-art; see A. If we consider high accuracy, then we can select SpikeNAS_2C2O_M with 92.02% accuracy and 19% memory saving from the state-of-the-art (i.e., 14% memory saving from the given budget). For CIFAR100, SNASNet incurs 23% memory overhead than the given budget, while SpikeNAS can save the memory footprint up to 71% from the state-of-the-art; see B. If we consider high accuracy, then we can select SpikeNAS_2C2O_M with 68.41% accuracy and 32% memory saving from the state-of-the-art (11% memory saving from the given budget). Meanwhile, for TinyImageNet200, SNASNet and our SpikeNAS meet the given memory budget. Moreover, SpikeNAS can save the memory footprint up to 71% from the state-of-the-art; see C. If we consider high accuracy, then we can select SpikeNAS_2C3O_M with 72% accuracy and 0.4% memory saving from the given budget.
Furthermore, our quantization further reduces the memory footprint of SNN models. Fig. 12(b) presents the results of experimental case study that compares SNASNet and our SpikeNAS_2C2O_M under different precision levels. Here, we observe that the high accuracy of SpikeNAS_2C2O_M can be maintained as long as at least 10-bit precision is employed for CIFAR10 and CIFAR100, hence leading to 75% and 79% memory savings over SNASNet, respectively. Meanwhile for TinyImageNet200, the high accuracy of SpikeNAS_2C2O_M can be maintained as long as at least 8-bit precision is employed, hence leading to 63% memory saving over SNASNet.
These results highlight that, SpikeNAS can generate SNN architectures that meet the given memory constraints due to its memory-aware search algorithm, enabling efficient deployments of SNN models on the target hardware platforms.
5.3 Comparing Accuracy and Memory of NAS for SNNs
To show the advancements of NAS for SNNs in terms of accuracy and memory requirements, we compare our SpikeNAS against different state-of-the-art works (i.e., AutoSNN [5], AutoSpikformer [40], AutoST [46], and SNASNet [28]); see a summary in Table 3. If we consider memory constraints of 1.2M parameters for CIFAR10, 2M parameters for CIFAR100, and 4M parameters for TinyImageNet200, then our SpikeNAS achieves competitive accuracy as compared to other NAS methods, while meeting the given memory constraints. Meanwhile, other NAS methods achieve high accuracy but at the cost of significantly larger memory requirements (i.e., higher number of parameters) than our SpikeNAS, hence they fail to meet the given memory constraints. The reason is that, only our SpikeNAS framework has the focus on hardware resources in the NAS process, hence having the capabilities to meet the given memory constraints.
Table 3: Comparison of different state-of-the-art NAS methods for SNNs on CIFAR10, CIFAR100, and TinyImageNet200.
Note: results for [5, 40, 46] are obtained from the respective papers; results for [28] and our SpikeNAS are obtained from our experiments; N/A means no available information; while results from hardware-aware NAS are written in bold. (*) Only our SpikeNAS considers hardware resources in the NAS process.
Work CIFAR10 CIFAR100 TinyImageNet200
# Param.
[M]
Acc.
# Param.
[M]
Acc.
# Param.
[M]
Acc.
[5]
1.46
20.92
91.32
93.15
N/A 69.16 N/A 46.79
[40]
4.69
9.20
95.29
96.38
4.69
9.20
77.71
77.05
N/A N/A
[46]
11.52
29.64
96.03
96.21
11.52
29.64
79.44
79.69
5.99
34.44
63.80
74.54
[28] 1.28 92.58 2.59 69.18 2.59 53.00
Ours
1.04
7.38
92.02
93.63
1.77
7.67
68.41
73.74
3.98
5.90
72.00
66.00
5.4 Reducing the Searching Time
@@
Fig. 13 presents the experimental results of the searching time. In all cases of workloads (i.e., CIFAR10, CIFAR100, and TinyImageNet200), our SpikeNAS offers significant improvements for the searching time as compared to the state-of-the-art SNASNet. Specifically, our SpikeNAS expedites the searching time by 1.48x-695x for CIFAR10 G, by 1.88x-247x for CIFAR100 H, and by 2.06x-63x for TinyImageNet200 I. The reason is that, SpikeNAS employs fewer operation types that have significant impact on the accuracy, performs a consecutive search across the existing cells, and incorporates a memory constraint into the search process, thereby leading to a smaller search space and a faster searching time than the state-of-the-art. If we consider high accuracy, acceptable memory cost, and fast searching time, then we can select SpikeNAS_2C2O_M for CIFAR10 with 92.02% accuracy, 14% memory saving from the given memory budget, and 29x faster search than the state-of-the-art. For CIFAR100, we can select SpikeNAS_2C2O_M with 68.41% accuracy, 11% memory saving from the given memory budget, and 117x faster search than the state-of-the-art. Meanwhile for TinyImageNet200, we can select SpikeNAS_2C3O_M with 72% accuracy, 0.38% memory saving from the given memory budget, and 3.7x faster search than the state-of-the-art.
These results highlight that, our SpikeNAS can quickly generate SNN architectures that offer high accuracy and meet the memory constraints, hence enabling efficient design automation for developing SNN-based embedded AI systems.
Figure 13: Results of the searching time normalized to the state-of-the-art work SNASNet_2C5O for (a) CIFAR10, (b) CIFAR100, and (c) TinyImageNet200 across different NAS techniques and settings.
Figure 14: Results of the energy consumption normalized to the state-of-the-art work SNASNet_2C5O for SNN implementation on (a) Nvidia RTX A6000 GPU machine, and (b) Nvidia Jetson AGX Orin, across different NAS techniques, settings, and datasets: CIFAR10, CIFAR100, and TinyImageNet200.
5.5 Improving Energy Efficiency
Fig. 14 presents the experimental results of the energy consumption for SNN implementation on Nvidia RTX A6000 GPU machine (represents a desktop platform) and Nvidia Jetson AGX Orin (represents an embedded platform). In general, our SpikeNAS incurs significantly lower energy consumption as compared to the state-of-the-art SNASNet in both desktop and embedded platforms, therefore leading to notable energy efficiency improvements. Specifically, in the RTX A6000 GPU machine, our SpikeNAS reduces energy consumption by 33.07%-99.86% (i.e., 1.49x-711x energy efficiency improvements) for CIFAR10 J, by 32.57%-99.43% (i.e., 1.48x-17x energy efficiency improvements) for CIFAR100 K, and by 43.57%-98.95%x (i.e., 1.77x-95x energy efficiency improvements) for TinyImageNet200 L. Meanwhile, in the Jetson AGX Orin, our SpikeNAS reduces energy consumption by about 49%-99% (i.e., 2x-82x energy efficiency improvements) for CIFAR10 M, CIFAR100 N, and TinyImageNet200 O. These energy savings come from the effectiveness of our proposed optimization techniques in SpikeNAS that limit the network size based on the memory constraint, which lead to the reduction of processing power and processing time, and hence improving the energy efficiency.
5.6 Flexibility of SpikeNAS Framework
In the SNN community, many different optimization techniques have been explored to improve the energy efficiency of SNN processing [52, 53, 54], and they also have the potential to be incorporated into our SpikeNAS framework to further enhance the efficiency of SNN models. For instance, the SATA accelerator [52] can be employed to expedite the training of SNN architectures that are generated from SpikeNAS, since both SATA accelerator and SpikeNAS framework consider the same surrogate gradient learning [51] for SNN training. The MINT method [53] can be employed in the quantization step of SpikeNAS for providing more options of quantization modes and memory saving, since MINT mainly focuses on quantizing the membrane potentials of LIF neurons. Meanwhile, a workload-balanced pruning technique [54] can be performed after the training phase in SpikeNAS to further compress the SNN model while ensuring a balanced-workload during the inference phase. Therefore, SpikeNAS framework has high flexibility for integration with other optimization techniques.
Our proposed techniques in the SpikeNAS framework (e.g., reduction of network operations) could be applicable to ANNs. For instance, the work of [55] may reduce the number of operations through pruning to further improve its efficiency. However, SNN-based solutions can offer alternate and unique advantages over ANN-based solutions, such as their ultra-low power/energy computation due to sparse spike-based computation and their inherent spatio-temporal learning capabilities, which are especially important for resource-constrained embedded AI systems [56]. To enable the SNN applicability for diverse application use-cases, SNN developments should be performed considering the design requirements (e.g., memory constraint), and supported by the bio-inspired computing platform with minimum/no power-hungry hardware like Analog-to-Digital Converters (ADCs) [56][57]. Therefore, our SpikeNAS framework enables a fast memory-aware NAS in developing SNN model that meets the given design requirements. Since NAS techniques for SNNs have not been extensively explored, our SpikeNAS framework may pave a way in enabling efficient SNN design targeting resource-constrained embedded AI systems.
6 Conclusion